# Stabilizing and Solving Inverse Problems using Data and Machine Learning

### v2

@@ 4.3 Inverse Problem
(b) Operator network $\phi_{u,N,h}$
@@ 5.1 Linear Operator with Linear Data
We use the truncated POD boundary basis corresponding to significant singular values to compute and interior basis. We do this by solving Laplace’s equation with the finite element method (FEM). We take the discrete space $V_{h}$ to simply be the space of piecewise linear finite elements on the triangle mesh considered. The FEM interior basis $\{\varphi_{\mathrm{FEM},n}\}_{n=1}^{N}$ is computed by: For each $n=1,...,N$, find $\varphi_{\mathrm{FEM},n}\in V_{h}$ such that $\varphi_{\mathrm{FEM},n}|_{\partial\Omega}=\varphi_{\mathrm{POD},n}$ and
@@ 5.2 Nonlinear Operator with Linear Data
This functional corresponds to the nonlinear differential operator whose inverse (the solution operator) we want to approximate with the neural network. We use a simple multilayer perceptron (MLP) network architecture with 4 hidden layers of the same width X and an output layer of width O representing the finite element DoFs. For standard P1 elements considered here it is simply the finite element function’s nodal values. We use the exponential linear unit (ELU) as the activation function in the 4 hidden layers and no activation function in the last layer. A schematic illustration of this network is provided in Figure 2(b).
@@
The hyperparameter values for the problem sizes have been obtained by trial and error. In Table 2, we present training info for the four mesh sizes for coefficient vector length $N=9$ and $N=21$. The training has been performed on a single A100 GPU. For the largest mesh case (244x244), we have not been able to train with all elements present in the energy functional loss function (It has resulted in a NaN loss function value). To make it work, we have employed the trick of randomly selecting a fixed number of elements for every input vector during training, and only considering the energy functional contribution from those elements. The number of elements used is denoted “Els” in Table 2. It can be observed from the zero energies and norm errors in Table 2 that the operator networks generally become more accurate with finer meshes if all elements are used in the energy computation. In both cases with the 244x244 mesh, the trend in higher accuracy is broken. This is reasonable since only a few elements, instead of all, are used in the energy computation for the loss function.
@@ 5.3 Nonlinear Operator with Nonlinear Data
We analyze data generated for the linear case with PCA and data generated for the quadratic case with both PCA and autoencoders. For the autoencoders we have used MLPs with 6 layers (5 hidden, 1 output) with the third layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 240 – 270 s.
@@
We analyze data generated for these cases with both PCA and autoencoders. For the autoencoders we have used MLPs with 5 layers (4 hidden, 1 output) with the middle layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 210 – 250 s.
@@
Table 3: Summary of optimization results for all nonlinear inverse problems using an operator network. All problems have the same reference solution and use the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs). The optimization processes for the problems are presented in Figures 6, 1012, respectively. In the table, “Op” means the operator network, “dec” means decoder, “sd = x” means what perturbation was added to the training data for the decoder, and “$\delta_{\omega}$” means that noisy data was used for the inverse problem.

### v3

@@ 4.3 Inverse Problem
(b) Operator network $\phi_{u,N,h}$ with four nonlinear layers (purple) and one linear output layer (blue)
@@ 5.1 Linear Operator with Linear Data
We use the truncated POD boundary basis corresponding to significant singular values to compute and interior basis. We do this by solving Laplace’s equation with FEM. We take the discrete space $V_{h}$ to simply be the space of piecewise linear finite elements on the triangle mesh considered. The FEM interior basis $\{\varphi_{\mathrm{FEM},n}\}_{n=1}^{N}$ is computed by: For each $n=1,...,N$, find $\varphi_{\mathrm{FEM},n}\in V_{h}$ such that $\varphi_{\mathrm{FEM},n}|_{\partial\Omega}=\varphi_{\mathrm{POD},n}$ and
@@ 5.2 Nonlinear Operator with Linear Data
This functional corresponds to the nonlinear differential operator whose inverse (the solution operator) we want to approximate with the neural network. We use a simple multilayer perceptron network architecture with 4 hidden layers of the same width X and an output layer of width O representing the finite element DoFs. For standard P1 elements considered here it is simply the finite element function’s nodal values. We use the exponential linear unit (ELU) as the activation function in the 4 hidden layers and no activation function in the last layer. A schematic illustration of this network is provided in Figure 3(b).
@@
The hyperparameters of the network (number of layers, width, activation function) and training settings (optimizer, number of iterations, decreasing learning rate, batch size, etc.) have been obtained by trial and error, where we have looked at the zero energy. An intuition for the size of the network (number of layers and widths) is that it needs to be large enough to provide a good approximation of the solution operator, but sufficiently small so that training is feasible and economical. An intuition for using ELU is that it is smoother than many other activations, e.g., the commonly used ReLU, which is to some degree in accordance with PDE- theory, where the solution is expected to depend smoothly on the problem data, e.g., boundary values.
In Table 2, we present training info for the four mesh sizes for coefficient vector length $N=9$ and $N=21$. The training has been performed on a single A100 GPU. For the largest mesh case (244x244), we have not been able to train with all elements present in the energy functional loss function (It has resulted in a NaN loss function value). To make it work, we have employed the trick of randomly selecting a fixed number of elements for every input vector during training, and only considering the energy functional contribution from those elements. The number of elements used is denoted “Els” in Table 2. It can be observed from the zero energies and norm errors in Table 2 that the operator networks generally become more accurate with finer meshes if all elements are used in the energy computation. In both cases with the 244x244 mesh, the trend in higher accuracy is broken. This is reasonable since only a few elements, instead of all, are used in the energy computation for the loss function.
@@ 5.3 Nonlinear Operator with Nonlinear Data
We analyze data generated for the linear case with PCA and data generated for the quadratic case with both PCA and autoencoders. For the autoencoders we have used MLPs with 6 layers (5 hidden, 1 output) with the third layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. The hyperparameters of the autoencoders and training settings have, just as in the case of the operator networks, been obtained by trial and error. To measure well-trainedness, we have looked at the average mean squared reconstruction error over 1000 unperturbed samples generated in the same way as during training. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 240 – 270 s.
@@
We analyze data generated for these cases with both PCA and autoencoders. For the autoencoders we have used MLPs with 5 layers (4 hidden, 1 output) with the middle layer being the latent layer. The latent layer width has been varied and the remaining hidden layer widths have all been fixed at 64. The activation function ELU has been applied to all layers except the last. The training has been performed with the Adam optimizer exactly as for the operator networks, i.e., $10^{6}$ iterations with a decreasing learning rate. The batch size has been 64. Again, the hyperparameters of the autoencoders and training settings have been obtained by trial and error, where we have looked at the average mean squared reconstruction error over 1000 unperturbed samples generated in the same way as during training. The training time for a single autoencoder (fixed latent layer width) on an Apple M1 CPU has typically been in the range 210 – 250 s.
@@
Table 3: Summary of optimization results for all nonlinear inverse problems using an operator network. All problems have the same reference solution and use the operator network in the last row of Table 2(b) (21 input coefficients, 59049 output DoFs). The optimization processes for the problems are presented in Figures 7, 1113, respectively. In the table, “Op” means the operator network, “dec” means decoder, “sd = x” means what perturbation was added to the training data for the decoder, and “$\delta_{\omega}$” means that noisy data was used for the inverse problem.